Stefani Company has gathered the following information about its product.

Direct materials: Each unit of product contains 3.10 pounds of materials. The average waste and spoilage per unit produced under normal conditions is 0.90 pounds. Materials cost $4 per pound, but Stefani always takes the 1.00% cash discount all of its suppliers offer. Freight costs average $0.25 per pound.

Direct labor. Each unit requires 2.10 hours of labor. Setup, cleanup, and downtime average 0.20 hours per unit. The average hourly pay rate of Stefani’s employees is $12.60. Payroll taxes and fringe benefits are an additional $2.30 per hour.

Manufacturing overhead. Overhead is applied at a rate of $4.90 per direct labor hour.

Compute Stefani’s total standard cost per unit. (Round answer to 2 decimal places, e.g. 1.25.)

Total standard cost per unit $
Answer #1
Answer
Material $ 16.84
Labour $ 34.27
Overhead $ 11.27
Total std. cost per unit = $ 62.38
Working
Material
Requirements = 3.10 + 0.90 ( wastage ) = 4 pounds
Cost = $ 4 * 99 % = $ 3.96
Freight = 0.25 * 4 = 1
Total = 3.96 * 4 + 1= $ 16.84
Labour
Requirement = 2.1 hour + 0.20 ( downtime ) = 2.30 hr.
Rate = $ 12.6
Payroll =2.30 * 2.30
Total = 2.30 * 12.6 + 2.30 * 2.30 = $ 34.27
Overhead
4.9 * 2.3 = $ 11.27